# Are Time-Indexed Foundation Models the Future of Time Series Imputation?

### v1

@@ B.1 Univariate datasets
The complete list of datasets used in our univariate experiments is shown in Table˜3. 36 datasets were used in total: 3 for the pretraining of MoTM and the remaining 33 for the zero-shot evaluation of both TabPFN-TS and MoTM. Each dataset is split chronologically intro train, validation and test. Unless otherwise stated, the respective fractions are 0.7, 0.1 and 0.2. The test segments were then generated by applying a four-week sliding window, where at every step a random stride is drawn uniformly between 0.5 and 2 days. This procedure ensures that the inference samples are not aligned on any specific calendar information. A short description of each dataset is provided below. Those datasets marked with an "*" were curated to remove flat segments at inference to avoid biasing the evaluation towards trivial scenarios. Flat segments are caused e.g. by heterogeneous sensor operating dates within datasets or by filling long missing blocks with zeros (Emami et al., 2023).
@@ B.2 Datasets with covariates
Table˜4 shows the key properties of the three datasets used for the evaluation of imputation with additional covariates. Each dataset is split chronologically intro train, validation and test splits with respective fractions 0.7, 0.1 and 0.2. Four-week segments are generated in the same manner as in Section˜B.1 for the univariate experiments.
@@ C.1 Full results
Section˜C.1 reports the complete univariate imputation benchmark across all datasets and missingness settings. These detailed results complement the main text by providing per-dataset performance for every method, allowing a finer comparison of their robustness and consistency across different missingness patterns and time series domain. For more details about the datasets please refer to Section˜B.1. All experiments were carried out on single NVIDIA A100-40G or H100-80G GPUs.
@@ Results.
Table˜5 reports the WQL scores across all eleven datasets and four missingness settings. TabPFN-TS achieves the lowest WQL across most datasets, showing its ability to produce both accurate and well-calibrated uncertainty estimates. MoTM follows with higher WQL but consistent behavior across datasets and settings. CSDI ranks third overall, with performances illustrating its limited generalization abilities: (i) consistently with its training procedure, good scores are obtained in the pointwise settings; (ii) except on small-size training sets, such as Covid19 Energy or PDB; (iii) CSDIgenerally suffers from significantly higher losses on the unseen Block settings. Finally, the SAITS variant adapted for quantile prediction yields the highest WQL in most settings, highlighting the shortcomings of this straightforward per-quantile adaptation compared with models expressly designed for probabilistic imputation.
@@ C.4 Experiments on lower sampling rates datasets
We consider four publicly available low-frequency datasets: Births Daily, M4 Daily, Births Weekly, and M4 Weekly, with statistics summarized in Table˜6. They are part of the Monash Time Series Forecasting archive (Godahewa et al., 2021), and were downloaded from the GIFT-eval repository (Aksu et al., 2024). Each dataset exhibits distinct temporal behaviors—seasonality and periodicity are typically weaker at weekly scales, while daily data contain more regular cycles and higher variance. For each dataset, we apply the same four missingness regimes as in the main benchmark (two pointwise and two block-based scenarios), allowing a consistent comparison across frequencies.
@@
We evaluate a representative subset of imputation methods present in the main benchmark: local heuristics (Linear, Seasonal Naive), supervised models SAITS, CSDI, BRITS, and the two foundation models TabPFN-TS and MoTM. As before, TabPFN-TS and MoTM operate in a fully zero-shot setting, while SAITS is retrained for each dataset. All metrics are aggregated using the z-normalized Mean Absolute Error (MAE), and results are reported per dataset and missingness pattern in Table˜7.
@@ Results.
As shown in Table˜7, foundation models demonstrate robust performance on low-frequency data. TabPFN-TS achieves the best average score (0.285), followed by MoTM (0.351), confirming their ability to generalize to coarser temporal structures without retraining. Other baselines, including specialized models like SAITS and BRITS, are significantly outperformed and show no clear advantage over simpler methods, while CSDI struggles notably. Overall, these results highlight the superior generalization of our zero-shot models in low-frequency regimes compared to both classic baselines and other deep learning architectures.
Appendix D Experiments with covariates: extensive qualitative results
@@
To better visualize the impact of incorporating covariates, we present qualitative results for both TabPFN-TS and MoTM. Figure˜10 and Figure˜11 show examples from three datasets (PV-France, Wind-France, and Load-France), illustrating how the inclusion of covariate information helps each model reconstruct the four one-day missing blocks more accurately.
@@ Results.
As illustrated in Figure˜10 and 11, incorporating covariates visually improves the reconstructions for both TabPFN-TS and MoTM. The covariate-enhanced versions capture missing intervals more smoothly and better follow short-term temporal variations within the four one-day gaps. We observe that the interpolations of TabPFN-TS align more closely with the ground truth and preserve sharper transitions across missing regions. In contrast, if MoTM consistently benefits from covariate information, its reconstructions sometimes struggle to capture sudden changes with great accuracy. Overall, the qualitative plots confirm that the most visible gains are observed on the Wind-France and PV-France datasets, while the impact remains limited for Load-France. One might note that quantile predictions can also be generated in this covariate setting, in a similar fashion to the univariate examples presented in Figure˜11.

### v2

@@ B.1 Univariate datasets
The complete list of datasets used in our univariate experiments is shown in Table˜4. 36 datasets were used in total: 3 for the pretraining of MoTM and the remaining 33 for the zero-shot evaluation of both TabPFN-TS and MoTM. Each dataset is split chronologically intro train, validation and test. Unless otherwise stated, the respective fractions are 0.7, 0.1 and 0.2. The test segments were then generated by applying a four-week sliding window, where at every step a random stride is drawn uniformly between 0.5 and 2 days. This procedure ensures that the inference samples are not aligned on any specific calendar information. A short description of each dataset is provided below. Those datasets marked with an "*" were curated to remove flat segments at inference to avoid biasing the evaluation towards trivial scenarios. Flat segments are caused e.g. by heterogeneous sensor operating dates within datasets or by filling long missing blocks with zeros (Emami et al., 2023).
@@ B.2 Datasets with covariates
Table˜5 shows the key properties of the three datasets used for the evaluation of imputation with additional covariates. Each dataset is split chronologically intro train, validation and test splits with respective fractions 0.7, 0.1 and 0.2. Four-week segments are generated in the same manner as in Section˜B.1 for the univariate experiments.
@@ C.1 Full results
LABEL:tab:full-results reports the complete univariate imputation benchmark across all datasets and missingness settings. These detailed results complement the main text by providing per-dataset performance for every method, allowing a finer comparison of their robustness and consistency across different missingness patterns and time series domain. For more details about the datasets please refer to Section˜B.1. All experiments were carried out on single NVIDIA A100-40G or H100-80G GPUs.
@@ Results.
Table˜6 reports the WQL scores across all eleven datasets and four missingness settings. TabPFN-TS achieves the lowest WQL across most datasets, showing its ability to produce both accurate and well-calibrated uncertainty estimates. MoTM follows with higher WQL but consistent behavior across datasets and settings. CSDI ranks third overall, with performances illustrating its limited generalization abilities: (i) consistently with its training procedure, good scores are obtained in the pointwise settings; (ii) except on small-size training sets, such as Covid19 Energy or PDB; (iii) CSDIgenerally suffers from significantly higher losses on the unseen Block settings. Finally, the SAITS variant adapted for quantile prediction yields the highest WQL in most settings, highlighting the shortcomings of this straightforward per-quantile adaptation compared with models expressly designed for probabilistic imputation.
@@ C.4 Experiments on lower sampling rates datasets
We consider four publicly available low-frequency datasets: Births Daily, M4 Daily, Births Weekly, and M4 Weekly, with statistics summarized in Table˜7. They are part of the Monash Time Series Forecasting archive (Godahewa et al., 2021), and were downloaded from the GIFT-eval repository (Aksu et al., 2024). Each dataset exhibits distinct temporal behaviors—seasonality and periodicity are typically weaker at weekly scales, while daily data contain more regular cycles and higher variance. For each dataset, we apply the same four missingness regimes as in the main benchmark (two pointwise and two block-based scenarios), allowing a consistent comparison across frequencies.
@@
We evaluate a representative subset of imputation methods present in the main benchmark: local heuristics (Linear, Seasonal Naive), supervised models SAITS, CSDI, BRITS, and the two foundation models TabPFN-TS and MoTM. As before, TabPFN-TS and MoTM operate in a fully zero-shot setting, while SAITS is retrained for each dataset. All metrics are aggregated using the z-normalized Mean Absolute Error (MAE), and results are reported per dataset and missingness pattern in Table˜8.
@@ Results.
As shown in Table˜8, our foundation models demonstrate robust performance on low-frequency data. TabPFN-TS achieves the best average score (0.285), followed by MoTM (0.351), confirming their ability to generalize to coarser temporal structures without retraining. Other baselines, including specialized models like SAITS and BRITS, are significantly outperformed and show no clear advantage over simpler methods, while CSDI struggles notably. Overall, these results highlight the superior generalization of our zero-shot models in low-frequency regimes compared to both classic baselines and other deep learning architectures.
Appendix D TabPFN-TS for imputation: additional analysis
D.1 TabPFN-TS sensitivity to feature design
As shown in Figs.˜1 and 2, TabPFN-TS delivers remarkably strong zero-shot imputation performance, both in the purely univariate setting and when auxiliary covariates are provided. However, one limitation of the current formulation is that the time-index features fed to the model are manually designed (Cai et al., 2025). This raises an important question: to what extent does TabPFN-TS rely on these specific features, and how sensitive is its performance to alternative time encodings? This section investigates this question by systematically varying the feature set and quantifying its impact on imputation quality.
Setting.
We evaluate five alternative feature configurations across 11 representative datasets and the four missingness scenarios used throughout the benchmark. Each configuration corresponds to a different encoding of the time index:
  • TabPFN1: the original feature set used in the paper, combining the raw grid-normalized time index with daily and weekly sinusoidal encodings:
    $H(t)=\left(t,\;\sin\!\left(\tfrac{2\pi t}{P_{\text{day}}}\right),\;\cos\!\left(\tfrac{2\pi t}{P_{\text{day}}}\right),\;\sin\!\left(\tfrac{2\pi t}{P_{\text{week}}}\right),\;\cos\!\left(\tfrac{2\pi t}{P_{\text{week}}}\right)\right).$
  • TabPFN2: daily-only sinusoidal encodings:
    $H(t)=\left(\sin\!\left(\tfrac{2\pi t}{P_{\text{day}}}\right),\;\cos\!\left(\tfrac{2\pi t}{P_{\text{day}}}\right)\right).$
  • TabPFN3: raw grid-normalized time index only:
    $H(t)=(t).$
  • TabPFN4: purely random periodic features using random periods $P_{\mathrm{rand}_{1}}$ and $P_{\mathrm{rand}_{2}}$:
    $H(t)=\left(\sin\!\left(\tfrac{2\pi t}{P_{\mathrm{rand}_{1}}}\right),\;\cos\!\left(\tfrac{2\pi t}{P_{\mathrm{rand}_{1}}}\right),\;\sin\!\left(\tfrac{2\pi t}{P_{\mathrm{rand}_{2}}}\right),\;\cos\!\left(\tfrac{2\pi t}{P_{\mathrm{rand}_{2}}}\right)\right).$
  • TabPFN5: the original features augmented with the random features used in TabPFN4:
    $H(t)=\big(t,\;\sin(\tfrac{2\pi t}{P_{\text{day}}}),\;\cos(\tfrac{2\pi t}{P_{\text{day}}}),\;\sin(\tfrac{2\pi t}{P_{\text{week}}}),\;\cos(\tfrac{2\pi t}{P_{\text{week}}}),\;\sin(\tfrac{2\pi t}{P_{\mathrm{rand}_{1}}}),\;\cos(\tfrac{2\pi t}{P_{\mathrm{rand}_{1}}}),\;\sin(\tfrac{2\pi t}{P_{\mathrm{rand}_{2}}}),\;\cos(\tfrac{2\pi t}{P_{\mathrm{rand}_{2}}})\big)$
Each variant probes a distinct aspect of the temporal encoding: (i) removing weekly structure and time index (TabPFN2) tests whether multi-frequency seasonality and the relative time index information are essential; (ii) removing the raw time index (TabPFN2, TabPFN4) evaluates whether relative positional information is necessary; (iii) retaining only the grid index (TabPFN3) probes how much the model can extrapolate without any periodic cues; (iv) replacing meaningful periodicities with random ones (TabPFN4) tests the robustness of learned priors to feature misspecification; (v) adding spurious features (TabPFN5) assesses whether irrelevant periodicities degrade performance by confusing the model’s inductive bias. Together, these experiments provide a systematic view of how much TabPFN-TS depends on feature engineering versus its own learned prior.
Setting.
These variants are evaluated on eleven univariate datasets, namely: BDG2-Bear & Rat, Covid19 Energy, GFC12 Load, Hog, Jena Weather 10T, Jena Weather 1H, Oikolab Weather, PDB, Pedestrian Counts and Weather, with details given in Appendix˜B. Table˜9 reports the complete MAE results across all eleven datasets and four missingness scenarios.
Table 9: Complete MAE results of five TabPFN-TS temporal encoding variants on 11 datasets. For each setting, the best score (lowest) is in bold and the second best is underlined.
Dataset Setting
TabPFN1
TabPFN2
TabPFN3
TabPFN4
TabPFN5
BDG2-Bear Pointwise 1 0.171 0.502 0.234 0.942 0.171
Pointwise 2 0.223 0.552 0.317 0.971 0.229
Blocks 1 0.272 0.498 0.763 0.874 0.287
Blocks 2 0.280 0.492 0.779 0.899 0.296
BDG2-Rat Pointwise 1 0.196 0.921 0.261 0.539 0.198
Pointwise 2 0.256 0.607 0.332 0.956 0.263
Blocks 1 0.349 0.528 0.680 0.872 0.366
Blocks 2 0.355 0.524 0.690 0.895 0.366
Covid19 Energy Pointwise 1 0.075 0.397 0.171 0.969 0.078
Pointwise 2 0.132 0.454 0.284 1.004 0.134
Blocks 1 0.202 0.363 0.774 0.909 0.203
Blocks 2 0.201 0.365 0.784 0.936 0.204
GFC12 Load Pointwise 1 0.143 0.537 0.280 0.935 0.144
Pointwise 2 0.237 0.615 0.386 0.969 0.237
Blocks 1 0.353 0.521 0.703 0.864 0.354
Blocks 2 0.363 0.519 0.719 0.903 0.364
Hog Pointwise 1 0.196 0.635 0.230 0.879 0.196
Pointwise 2 0.260 0.716 0.287 0.924 0.261
Blocks 1 0.396 0.626 0.532 0.863 0.397
Blocks 2 0.406 0.618 0.546 0.878 0.406
Jena Weather 10T Pointwise 1 0.086 0.629 0.132 0.710 0.086
Pointwise 2 0.101 0.636 0.135 0.769 0.101
Blocks 1 0.378 0.653 0.521 0.661 0.379
Blocks 2 0.384 0.647 0.535 0.660 0.384
Jena Weather 1H Pointwise 1 0.156 0.546 0.210 0.848 0.156
Pointwise 2 0.214 0.628 0.260 0.906 0.214
Blocks 1 0.366 0.544 0.517 0.791 0.367
Blocks 2 0.370 0.538 0.525 0.821 0.371
Oikolab Weather Pointwise 1 0.150 0.684 0.204 0.926 0.176
Pointwise 2 0.228 0.765 0.263 0.967 0.228
Blocks 1 0.449 0.681 0.605 0.886 0.450
Blocks 2 0.454 0.683 0.611 0.909 0.454
PDB Pointwise 1 0.062 0.345 0.207 1.020 0.063
Pointwise 2 0.119 0.387 0.324 1.033 0.119
Blocks 1 0.171 0.319 0.831 0.895 0.171
Blocks 2 0.177 0.332 0.855 0.947 0.177
Pedestrian Counts Pointwise 1 0.150 0.337 0.363 0.937 0.153
Pointwise 2 0.200 0.369 0.443 0.970 0.214
Blocks 1 0.172 0.327 0.863 0.820 0.174
Blocks 2 0.178 0.327 0.880 0.846 0.181
Weather Pointwise 1 0.257 0.655 0.273 0.915 0.259
Pointwise 2 0.311 0.739 0.325 0.952 0.316
Blocks 1 0.456 0.660 0.579 0.867 0.459
Blocks 2 0.455 0.656 0.584 0.889 0.455
Mean (Std) 0.252 (0.114) 0.554 (0.171) 0.415 (0.244) 0.904 (0.102) 0.260 (0.116)
Results.
Table˜9 shows a clear and consistent trend: the original feature set (TabPFN1) remains the strongest across virtually all datasets and settings. (i) Daily-only features (TabPFN2) perform substantially worse, highlighting the importance of capturing both daily and weekly periodicities. (ii) Using only the raw index (TabPFN3) yields reasonable but clearly inferior performance, especially in block-missing scenarios where periodic cues are most helpful. (iii) Random features degrade performance drastically (TabPFN4), confirming that TabPFN-TS is not invariant to arbitrary feature choice. (iv) Finally, augmenting correct features with random ones (TabPFN5) results in little to no degradation, indicating that the model is generally robust to irrelevant or noisy features and does not easily overfit to spurious temporal structure.
Overall, these results underline the importance of well-designed time features and suggest that part of TabPFN-TS’s effectiveness stems from the alignment between its meta-training distribution and the provided temporal encodings.
Future work.
An important direction for future research is the automatic construction or selection of temporal features. Our experiments highlight that TabPFN-TS is highly sensitive to the structure of the time encodings it receives. Instead of relying on manually crafted sinusoidal components, one could learn feature representations jointly with the model, or derive them through a separate meta-learning (like MoTM) or feature-search mechanism. Such an approach would be particularly valuable when dealing with highly heterogeneous datasets, spanning very different temporal granularities (minutes, hours, days, months) and periodic structures. Automatically discovering the most relevant periodicities and positional encodings would likely make TabPFN-TS more robust, more general, and less dependent on domain-specific feature engineering.
D.2 Extended discussion on TabPFN-TS computational bottlenecks
A practical limitation of TabPFN-TS lies in its inference cost (see Section˜4.1). On an NVIDIA H100 GPU, a single forward pass over a chunk of 672 time steps requires approximately one second. Depending on the computational resources available and the frequency at which predictions must be updated, this can be either acceptable or prohibitive. For example, batch offline imputation or low-frequency forecasting pipelines can easily amortize this cost, whereas real-time applications or systems operating on thousands of concurrent time series may find the latency challenging. Reducing inference time—through model distillation, sparse attention, chunk parallelization, or specialized kernels—thus represents a key avenue to make TabPFN-TS more broadly deployable in production environments.
In this regard, the recently introduced TabPFN 2.5 (Grinsztajn et al., 2025) appears to be a promising solution, as it substantially accelerates in-context regression and could therefore mitigate the current inference bottlenecks of TabPFN-TS.
Appendix E Experiments with covariates: extensive qualitative results
E.1 Qualitative results
@@
To better visualize the impact of incorporating covariates, we present qualitative results for both TabPFN-TS and MoTM. Fig.˜10 and Fig.˜11 show examples from three datasets (PV-France, Wind-France, and Load-France), illustrating how the inclusion of covariate information helps each model reconstruct the four one-day missing blocks more accurately.
@@ Results.
As illustrated in Fig.˜10 and 11, incorporating covariates visually improves the reconstructions for both TabPFN-TS and MoTM. The covariate-enhanced versions capture missing intervals more smoothly and better follow short-term temporal variations within the four one-day gaps. We observe that the interpolations of TabPFN-TS align more closely with the ground truth and preserve sharper transitions across missing regions. In contrast, if MoTM consistently benefits from covariate information, its reconstructions sometimes struggle to capture sudden changes with great accuracy. Overall, the qualitative plots confirm that the most visible gains are observed on the Wind-France and PV-France datasets, while the impact remains limited for Load-France. One might note that quantile predictions can also be generated in this covariate setting, in a similar fashion to the univariate examples presented in Fig.˜11.
Coarse granularity interpolation.
We extend our qualitative analysis to a practical scenario of interest, namely interpolating sparse yet regular observations, e.g. at a multi-hourly time-step, to a finer time resolution, e.g. an hourly time-step. Fig.˜12 illustrates the behaviour of TabPFN-TS when interpolating the Wind-France and PV-France datasets from a six-hourly to an hourly sampling rate. On PV-France, we note that TabPFN-TS maintains shape consistency, although certain artifacts can be observed (typically non-zero values at night). These can be corrected by incorporating covariate information. On Wind-France, the model produces smooth univariate interpolations (blue line). Interestingly, adding the covariate is detrimental in this case: the in-context learning-based adaptation procedure of TabPFN on a sparse six-hour grid is mislead into spurious correlations by the stronger stochasticity of the covariate.
(a) Wind-France
(b) PV-France
Figure 12: TabPFN-TS qualitative results with and without covariates in coarse granularity scenarios: context observed at a six-hourly sampling rate, imputation on a dense hourly grid.
E.2 Discussion: integration of incomplete covariate data
The integration mechanism used in Section˜3.2 simply concatenates the covariate value $c(t)$ to the representation $H(t)$ before feeding it to the regression head $r_{\theta}(t)$. This assumes that the covariate is fully observed and temporally aligned with the target series for every timestamp $t$. When this assumption fails (for example, if the covariate contains missing values, is recorded asynchronously, or is sampled at a different frequency) naively integrating it into the model is not straightforward.
Below, we discuss practical strategies to handle such situations. These solutions are closer to implementation tricks than to principled methodological contributions, but they can be useful in practice.
Partially observed but aligned covariates.
When the covariate is sampled on the same time grid as the target series but contains missing values, a straightforward approach is to first impute the covariate and then apply the standard integration procedure. This imputation can be performed using simple linear interpolation or more sophisticated approaches such as TabPFN-TS (or MoTM for faster imputation).
A limitation of this strategy is that the imputation of the covariate does not leverage the target series, which may lead to suboptimal reconstructions. More advanced approaches could jointly model both the target and the covariate — for example using implicit neural representations to obtain functional embeddings independent of the original sampling grid — but such ideas remain open research directions.
Fully observed but misaligned covariates (asynchronous sensors or differing sampling rates).
If the covariate is complete but not recorded on the same timestamps as the target series, simple resampling methods such as linear interpolation can be used to align it with the target grid. More flexible approaches, such as querying TabPFN-TS or MoTM at arbitrary timestamps, can also be applied since these models operate in a continuous-time manner rather than imputing values through mask-based reconstruction (Tashiro et al., 2021; Du et al., 2023a). This makes them naturally suitable for harmonizing covariates recorded at heterogeneous frequencies.
Appendix F Comparison with hyperparameters-optimized SAITS
The experiments reported in the main paper rely on the hyperparameters recommended by each supervised baseline in their respective publications. While this ensures a fair and reproducible benchmark, it is also informative to compare TabPFN-TS and MoTM against a version of SAITS whose hyperparameters have been actively tuned. This comparison is particularly relevant given that hyperparameter optimization constitutes one of the main disadvantages of supervised approaches, in contrast with zero-shot methods.
Setting.
We perform an explicit hyperparameter grid search for SAITS, varying its two main hyperparameters highlighted as important in the original paper, namely the inner model dimension ($d_{model}$) and the number of Transformer block layers ($n_{layers}$). The search is carried out on eleven representative datasets used throughout the benchmark—BDG2-Bear, BDG2-Rat, Covid19 Energy, GFC12 Load, Hog, Jena Weather 10T, Jena Weather 1H, Oikolab Weather, PDB, Pedestrian Counts, and Weather. For the hyperparameter search, the train/validation/test split follows a 0.7/0.1/0.2 ratio along the temporal axis, consistent with all other univariate experiments in the paper. For each dataset, we retain the hyperparameter configuration that yields the lowest validation MAE, averaged over the four validation-mask scenarios, and subsequently report the corresponding test MAE for each scenario. In addition, Fig.˜13 reports how the test MAE and training time vary with $d_{model}$ and $n_{layers}$ for SAITS.
Table 10: Best SAITS hyperparameter configuration per dataset based on validation MAE (averaged over the four validation-mask scenarios).
Dataset $d_{model}$ $n_{layers}$ Avg. Val. MAE
BDG2-Bear 128 4 0.276
BDG2-Rat 128 3 0.328
Covid19 Energy 128 3 0.213
GFC12 Load 128 3 0.289
Hog 128 4 0.416
Jena Weather 10T 256 2 0.367
Jena Weather 1H 128 4 0.390
Oikolab Weather 128 3 0.430
PDB 256 3 0.249
Pedestrian Counts 256 4 0.165
Weather 256 2 0.412
Table 11: Comparison of test MAE between the two zero-shot approaches, TabPFN-TS and MoTM, and the tuned SAITS baseline (best configuration selected per dataset via validation MAE). For each setting, the best (lowest) MAE is shown in bold, and the second best is underlined.
Dataset Setting
TabPFN-TS
MoTM
$\text{SAITS}_{opt}$
$\text{SAITS}_{base}$
BDG2–Bear Pointwise 1 0.171 0.202 0.191 0.241
Pointwise 2 0.223 0.240 0.253 0.309
Blocks 1 0.272 0.332 0.327 0.399
Blocks 2 0.280 0.336 0.325 0.405
BDG2–Rat Pointwise 1 0.196 0.231 0.199 0.266
Pointwise 2 0.256 0.273 0.270 0.339
Blocks 1 0.349 0.400 0.421 0.495
Blocks 2 0.355 0.402 0.420 0.497
Covid19 Energy Pointwise 1 0.075 0.099 0.344 0.399
Pointwise 2 0.132 0.127 0.380 0.417
Blocks 1 0.202 0.222 0.372 0.432
Blocks 2 0.201 0.232 0.367 0.436
GFC12 Load Pointwise 1 0.143 0.189 0.189 0.188
Pointwise 2 0.237 0.231 0.277 0.280
Blocks 1 0.353 0.382 0.399 0.417
Blocks 2 0.363 0.387 0.406 0.425
Hog Pointwise 1 0.196 0.240 0.226 0.245
Pointwise 2 0.260 0.286 0.292 0.320
Blocks 1 0.396 0.458 0.514 0.518
Blocks 2 0.406 0.464 0.516 0.528
Jena Weather 10T Pointwise 1 0.086 0.190 0.241 0.177
Pointwise 2 0.101 0.207 0.293 0.196
Blocks 1 0.378 0.453 0.502 0.557
Blocks 2 0.384 0.459 0.504 0.561
Jena Weather 1H Pointwise 1 0.156 0.224 0.241 0.346
Pointwise 2 0.214 0.272 0.293 0.394
Blocks 1 0.366 0.463 0.502 0.532
Blocks 2 0.370 0.464 0.504 0.532
Oikolab Weather Pointwise 1 0.150 0.227 0.188 0.207
Pointwise 2 0.228 0.278 0.293 0.294
Blocks 1 0.449 0.529 0.598 0.581
Blocks 2 0.454 0.534 0.602 0.584
PDB Pointwise 1 0.062 0.094 0.201 0.337
Pointwise 2 0.119 0.121 0.271 0.373
Blocks 1 0.171 0.197 0.317 0.353
Blocks 2 0.177 0.201 0.314 0.373
Pedestrian Counts Pointwise 1 0.150 0.196 0.128 0.240
Pointwise 2 0.200 0.239 0.183 0.289
Blocks 1 0.172 0.254 0.173 0.243
Blocks 2 0.178 0.260 0.176 0.252
Weather Pointwise 1 0.257 0.298 0.281 0.290
Pointwise 2 0.311 0.342 0.348 0.364
Blocks 1 0.456 0.494 0.556 0.592
Blocks 2 0.455 0.488 0.553 0.590
Mean (Std) 0.252 (0.114) 0.300 (0.122) 0.340 (0.159) 0.379 (0.122)
Results.
Table˜11 reports the detailed test MAE for each dataset and masking scenario. $\text{SAITS}_{opts}$ correspond to the best configurtaion and $\text{SAITS}_{base}$ correspond to the default hyperparameters baseline provided by the authors (see Section˜2.1). Table˜10 report the best configuration for each datasets according to the lowest average MAE obtained on the validation set. Overall, hyperparameter tuning yields only modest gains for SAITS: the tuned configuration slightly improves over the baseline, but is still outperformed, on average across the 11 datasets, by the two zero-shot approaches TabPFN-TS and MoTM by an average of 26% and 12% respectively. In addition, Fig.˜13(a) shows that, for $d_{model}\in\{128,256\}$ and $n_{layers}\in\{2,3,4\}$, the average test MAE remains essentially unchanged, indicating limited benefits from increasing depth within this range. By contrast, Fig.˜13(b) highlights that increasing either $d_{model}$ or $n_{layers}$ leads to a substantial increase in training time. These findings corroborate the default configuration proposed by the authors of SAITS (Du et al., 2023a), namely $d_{model}=256$ and $n_{layers}=2$.
Figure 13: Hyperparameter ablation for SAITS. (a) Average test MAE across the 11 datasets as a function of the number of layers, with one curve per $d_{model}$. (b) Average training time across the 11 datasets as a function of the number of layers, again with one curve per $d_{model}$.